Circular No. 8117

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003dm, 2003dn, 2003do, 2003dp, 2003dq, 2003dr Several additional supernovae found on unfiltered CCD images have

been reported — SN 2003dr in NGC 5714 independently by T. Puckett and D. Toth (cf. IAUC 8045) and by M. Schwartz and P. R. Holvorcem (LOTOSS/Tenagra II; cf. IAUC 8111), and the rest by W. M. Wood-Vasey, G. Aldering, P. Nugent, and A. Kulkarni (NEAT; cf. IAUC 8115). The tabulated data for SN 2003dr below are from Puckett.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003dm Mar. 22.20 11^{h}31^{m}38.^{s}15 +25^{o}34′14′′.3 19.2 2′′ W, 1′′ N
2003dn Apr. 7.30 11 20 08.03 +25 36 48.7 20.0 6′′ S
2003do Apr. 8.29 13 03 29.37 +27 19 11.1 19.9 1′′.4 E, 1′′.3 S
2003dp Apr. 8.36 13 13 24.23 +27 56 09.3 20.0 3′′.1 W
2003dq Apr. 10.15 11 04 45.44 +15 26 49.4 18.9 0′′.3 E, 1′′.9 S
2003dr Apr. 13.18 14 38 11.13 +46 38 03.4 18.0 3′′.7 W, 13′′.9 S

Additional approximate magnitudes from the respective observers: 2003dm, 2002 Mar. 31.38, [20.9; 2003 Jan. 19.54, [20.5; Apr. 7.33, 19.7. 2003dn, 2002 Mar. 31.37, [21.0; 2003 Feb. 21.50, [20.7; Apr. 10.15, 20.0. 2003do, 2002 Apr. 9.21, [21.6; 20.25, [20.9; 2003 Mar. 22.40, 20.2; Apr. 11.32, 19.9. 2003dp, 2002 May 3.21, [21.4; 2003 Mar. 22.40, [21.1; Apr. 11.32, 20.0. 2003dq, 2002 Apr. 2.27, [21.8; 2003 Mar. 24.16, [21.5; 25.24, [21.3; Apr. 9.42, 19.0. SN 2003dr, 1997 July 1, [21.0 (Palomar Sky Survey, via Puckett); 2003 Mar. 24, [20.5 (Puckett); Apr. 8.3, [18.5 (Holvorcem); 12.4, 17.9 (Holvorcem); 14.16, 18.0 (Puckett); 14.4, 17.9 (Holvorcem).

SUPERNOVAE 2003cx AND 2003dc M. Hamuy, Carnegie Observatories, reports that a spectrum (range

370–900 nm) of SN 2003cx (cf. IAUC 8105), obtained on Apr. 10.39 UT with the Las Campanas Observatory Clay 6.5-m telescope (+ LDSS-2), shows it to be a normal type-II supernova with Balmer lines exhibiting P-Cyg profiles. The expansion velocity of −10000 km/s, derived from the minimum of Hα (based on a redshift of 0.037, derived for the host galaxy from superimposed narrow emission Balmer lines), and the relatively blue continuum suggest that SN 2003cx is ∼ 2–3 weeks old. A spectrum of SN 2003dc (cf. IAUC 8111), taken on Apr. 10.37, shows it to be a type-Ia supernova near maximum light. The minimum of the Si II 635.5-nm line yields an expansion velocity of −10060 km/s (based on a redshift of 0.067, derived for the host galaxy from superimposed narrow Hα emission).

2003 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green